# Safeguarding Mobile GUI Agent via Logic-based Action Verification

### v1

@@ Abstract.
Large Foundation Models (LFMs) have unlocked new possibilities in human-computer interaction, particularly with the rise of mobile Graphical User Interface (GUI) Agents capable of interpreting GUIs. These agents promise to revolutionize mobile computing by allowing users to automate complex mobile tasks through simple natural language instructions. However, the inherent probabilistic nature of LFMs, coupled with the ambiguity and context-dependence of mobile tasks, makes LFM-based automation unreliable and prone to errors. To address this critical challenge, we introduce VeriSafe Agent (VSA): a formal verification system that serves as a logically grounded safeguard for Mobile GUI Agents. VSA is designed to deterministically ensure that an agent’s actions strictly align with user intent before conducting an action. At its core, VSA introduces a novel autoformalization technique that translates natural language user instructions into a formally verifiable specification, expressed in our domain-specific language (DSL). This enables runtime, rule-based verification, allowing VSA to detect and prevent erroneous actions executing an action, either by providing corrective feedback or halting unsafe behavior. To the best of our knowledge, VSA is the first attempt to bring the rigor of formal verification to GUI agent, effectively bridging the gap between LFM-driven automation and formal software verification. We implement VSA using off-the-shelf LLM services (GPT-4o) and evaluate its performance on 300 user instructions across 18 widely used mobile apps. The results demonstrate that VSA achieves 94.3%–98.33% accuracy in verifying agent actions, representing a significant 20.4%–25.6% improvement over existing LLM-based verification methods, and consequently increases the GUI agent’s task completion rate by 90%–130%.
copyright: noneconference: ; ;
@@ 1. Introduction
However, unlike conventional software verification (Liu et al., 2023; Yang et al., 2024; Wu et al., 2024; Obi et al., 2025), which operates on predefined safety requirements, VSA faces the unique challenge of formalizing impromptu, user-defined instructions on-the-fly, across diverse and dynamic landscape of mobile applications. To accomplish this, VSA introduces three key innovations:
@@
    Domain-Specific Language (DSL) and Developer Library: A specialized language and accompanying tools tailored to dynamic nature of mobile environments. Collectively, they can encode both the natural-language user instructions and corresponding UI actions as logical formulas.
@@
We implement a prototype of VSA using off-the-shelf online LLM service (GPT-4o) and integrate it with M3A (Multimodal Autonomous Agent for Android) GUI agent (Rawles et al., 2024). Our evaluation, conducted on 300 user instructions across 18 widely used mobile applications demonstrates that with an estimated 10 additional lines of code (LoC) per app, VSA successfully verifies GUI agent actions with up to 98.33% accuracy, achieving a false positive rate of 2.7% and a false negative rate of 0.7%. This significantly outperforms baseline reflection agents by 25.6%. Furthermore, VSA’s structured feedback, derived from logical verification results, enables the GUI agent to enhance its task completion rate by 130%, a substantial improvement compared to reflection agents, which failed to correct any tasks.
@@ 3.1. Challenges
C3. Simply flagging an action as incorrect is insufficient for improving the overall practicality of Mobile GUI Agents. To enhance their accuracy and reliability, we must provide actionable feedback that correctly guides agents toward the goal. However, LLM-generated feedback is often vague, underspecified, or even incorrect, potentially leading to repeated errors or suboptimal actions. Therefore, VSA incorporates a structured feedback generation mechanism (Figure 4) that systematically generates precise, rule-based feedback by identifying unmet or violated conditions based on the logical verification results.
@@ 3.2. System Workflow
App Development Phase. To enable rigorous rule-based verification, VSA requires app developers to define the critical states and transitions relevant for verification using VSA developer library (§ 4.2). This includes declaring application state spaces and their update mechanisms, which are already common practice in modern app development workflows (Meta, 2025; Google, 2025; Apple, 2025) and requires minimal additional effort, typically involving 5 to 10 lines of code per state.
@@
[…]
@@ 4.1. Design of Domain-Specific Language
Specification $::=\;\;\textit{Rule}^{*}$
[…]
Pred $::=\;\;\textit{Objective}\mid\textit{StatePred}(\textit{Constraint}^{*})$
[…]
@@
Constraint. A constraint $c$ has the form $var\;op\;const$ (e.g., $x=10$), where $x$ is a variable representing a specific element within the application state (e.g., restaurant_name, location). $op$ is a comparison operator, and $const$ represents a constant value. The expression $x;op;const$ returns true if the variable $x$ meets the condition defined by the operator $op$ and the constant $const$. If $x$ has not been observed in the app, it defaults to undefined, making the constraint always false. The variable $x$ can be one of the following six types: string, number, Boolean, enumeration, date, and time. For example, given a constraint $x\geq 10$, if the value of the variable $x$ was observed to be 15 in the application, then the constraint is considered true.
@@
[…]
@@
[…]
@@ 4.2. Developer Library
To this end, VSA provides a developer library that empowers app developers to explicitly define application states (as state predicates) and their corresponding update mechanisms (as state transitions). This approach capitalizes on developers’ deep understanding of their application’s internal logic and intended behavior. Although it requires additional developer effort, it enables highly accurate state definitions suitable for rule-based logical verification—a level of precision that is unattainable with automated approaches like static code analysis or LLM-driven dynamic analysis.
@@
safeMATE.updateState("RestaurantInfo", {
@@
Another key advantage of developer-defined state predicates is the control over granularity. In many mobile tasks, individual atomic actions are not inherently meaningful. For example, in an e-commerce application, entering individual fields of a shipping address (street, city, zip code) is a set of preparatory steps. The critical state change is the submission of the complete address. Verifying every single atomic action would be inefficient and could lead to false positives when verification occurs before completing all preparatory steps. By allowing developers to strategically define verification checkpoints that have a meaningful impact on the task’s execution, VSA can significantly improve the efficiency and accuracy of task verification.
@@ 5.1. Intent Encoder
To ensure the encoded specification is correct and includes all relevant constraints, VSA employs two complementary self-corrective techniques: (i) rule-based syntax checking, and (ii) decoding-based semantics checking. These methods allow the LLM to incrementally identify, correct, and refine its translations, improving accuracy and reliability.
As an initial step, the Intent Encoder prompts the LLM with: i) the user’s natural language instruction and ii) the set of developer-defined states (from the developer library). From this information and user instruction, the Intent Encoder generates a draft specification representing the user’s intent using available states. The draft specification then undergoes two checking stages:
Rule-based Syntax Checking. Inspired by conventional programming languages, VSA checks that the generated Specification strictly adheres to the DSL’s syntax (§ 4.1.1) and that the types of constants $c$ used within the Specification are consistent with the types of variables declared in the developer-defined states. For example, if the LLM generates a state predicate like $\textsf{RestaurantInfo}(\textsf{name}\geq\textsf{100})$ while the type of the name variable is defined as ‘String’ in the developer library, this inconsistency is immediately flagged as a syntax error. If an error is detected, a structured error message is returned to the encoding LLM as feedback, guiding it toward a corrected translation.
Decoding-Based Semantics Checking. To further validate the semantic correctness of the translation, the generated Specification is decoded back into a natural language description using a separate decoder LLM. A checker LLM then compares this decoded description to the original user instruction, confirming that the encoding genuinely captures the user’s intent rather than merely manipulating symbols. If discrepancies are found, the checker LLM identifies the root cause and provides feedback to the encoding LLM.
These complementary checking mechanisms significantly improve the robustness and accuracy of the intent encoding process, effectively mitigating risks associated with solely relying on probabilistic LLM outputs.
@@
To address this challenge, VSA introduces an Experience-driven Encoding, which leverages previously successful encodings to guide subsequent translations. When an instruction is successfully translated and verified, its result specification is cached. Since a single specification consists of multiple rules representing an instruction’s intermediate objectives, numerous instructions for the same application may share similar rules. By retrieving and prioritizing predicates based on past encoding experience, we can significantly reduce the search space for the LLM encoder, improving both efficiency and consistency.
@@ 5.2. Intent Verifier
[…]
@@
Rule-level Verification is performed at irreversible or critical checkpoints within the task execution. To identify these checkpoints, we modified the GUI agent’s prompt so that it generates a critical flag along with the action if that action is intended to directly achieve one of the objectives $o$ within the Specification (e.g., Reserve)
@@
These two-tiered verification approach guides the agent toward its objectives while ensuring the safe execution of high-risk actions.
@@ 7. Implementation
Our VSA implementation is designed as a modular component that can be seamlessly integrated with existing GUI agents without requiring modifications to their underlying architecture. VSA is implemented in using Python, and its Intent Encoder leverages off-the-shelf LFM GPT-4o.
Handling unpredictable state updates. While VSA focuses on pre-action verification, it also supports post-action updates where the outcome of an action is unpredictable or depends on external factors (e.g., the result of a network request). In these cases, developers can directly use the ‘updateState()’ API without a wrapper listener to update the state after the desired information is known. This allows for verification even in scenarios where pre-action prediction is impossible.
Developer Library Guidelines. Developers should carefully determine the granularity of State Predicates to suit their application context. Overly fine-grained State Predicates can make the encoded specification too detailed, increasing the false positive rate. Conversely, coarse-grained State Predicates may introduce loopholes in the encoded specification, leading to a higher false negative rate. Developers must find a sweet spot in this trade-off based on their application’s usage patterns. To assist with this, the VSA Library provides detailed documentation and guidelines on effectively defining and utilizing State Predicates to achieve robust verification. Given the potential severity of false negatives, it’s recommended to begin with a fine-grained verification first and then iteratively refine the predicates to minimize false positives.
String Similarity. From the perspective of structural equivalence, strings with different notations or synonymous words are treated as distinct entities (e.g., Jan 01 and January 1st). However, user instructions may not exactly match the values in the application, making it necessary to check for semantic equivalence rather than structural equivalence. To achieve this, we used OpenAI’s text-embedding-3-small model (openai, 2023) to generate semantic embeddings and compared them using cosine similarity. Based on empirical studies, we determined an equivalence threshold of 0.7.
@@ 8. Evaluation
Figure 5. Verification accuracy
In this section, we demonstrate the performance of VSA in three key aspects: i) verification accuracy, ii) feedback effect, and cost&latency. To this end, we deployed M3A (Rawles et al., 2024), a simple yet powerful mobile GUI agent powered by GPT-4o, on a Google Pixel 8 smartphone and integrated it with VSA. This setup allows VSA to validate the agent’s behavior in runtime while enabling autonomous app execution.
Baselines & VSA Variants. To evaluate VSA, we compare it with two LFM-based reflection schemes: Pre-action reflection (Li et al., 2024a) and Post-action reflection (Wang et al., 2024), both of which utilize GPT-4o as their base LFM. For VSA, we evaluate two variants: VSA-cold and VSA-warm. VSA-cold encodes user instructions solely using self-corrective encoding, without experience-driven encoding. On the other hand, VSA-warm leverages both self-corrective encoding and experience-driven encoding, assuming that a cached list of candidate predicates is available for the given instruction (see § 5.1 for details).
@@ 8.1. Dataset
The primary contribution of VSA is its ability to proactively verify and provide feedback to a GUI agent when automating mobile tasks. To comprehensively evaluate VSA’s effectiveness, we constructed a dataset consisting of 300 user instructions. This dataset is divided into two complementary subsets: Correct and Wrong datasets, each containing 150 instructions.
The Correct dataset represents scenarios where the GUI agent’s execution path precisely matches the user’s instruction. This set contains instructions paired with their ground-truth execution paths (sequences of UI actions that correctly fulfill the instruction). It is designed to evaluate True Negatives (correctly identifying a valid execution) and False Positives (incorrectly flagging a valid execution as erroneous). Within this dataset, we included 125 instructions from the widely adopted LlamaTouch dataset (Zhang et al., 2024), covering 18 mobile applications. However, since the LlamaTouch instructions are predominantly simple (average 5.67 steps), similar to other datasets (Wen et al., 2023; Rawles et al., 2024; Xing et al., 2024; Rawles et al., 2023), we augmented our dataset with 25 additional complex instructions, referred to as the Challenge dataset. These instructions feature instructions with significantly greater complexity, including longer action sequences, multiple constraints, and conditional branches. These challenging instructions span 9 different applications, with an average of 19.16 steps and a maximum length of 40 steps. An example of such a challenging instruction in Google Maps is: “I am planning to visit the Cattedrale di Santa Maria del Fiore soon. First, find the Cattedrale di Santa Maria del Fiore on the map and then look for nearby attractions. Next, identify one attraction with a rating of 4 stars or higher and at least 2000 reviews, then show me the wheelchair-accessible walking route from the Cattedrale di Santa Maria del Fiore to that attraction.”
To thoroughly evaluate error-detection performance, we constructed a complementary Wrong dataset by deliberately introducing mismatches between instructions and their corresponding execution paths from the Correct dataset. Specifically, we altered key details in the original instructions to create discrepancies. For instance, we modified the instruction: ”On Play Books, turn to the search page and check for top-selling books.” to: ”On Play Books, turn to the search page and check for newly released books,” Because the execution path no longer aligns with the modified instruction, these scenarios effectively test the verification system’s ability to detect errors— True Positive (correctly identified errors) and False Negative (missed errors).
Finally, as described in § 4.2, VSA relies on developer-defined state definitions and state-update triggers. However, since the LlamaTouch dataset uses commercial applications without available source code, we manually annotated each application’s GUI screens with the required predicate information and state transitions using a GUI annotation tool. This manual annotation achieves the same effect as direct integration via the developer library, effectively replicating the developer library’s intended functionality.

### v2

# VeriSafe Agent: Safeguarding Mobile GUI Agent via Logic-based Action Verification

@@ Abstract.
[…]
Large Foundation Models (LFMs) have unlocked new possibilities in human-computer interaction, particularly with the rise of mobile Graphical User Interface (GUI) Agents capable of interacting with mobile GUIs. These agents allow users to automate complex mobile tasks through simple natural language instructions. However, the inherent probabilistic nature of LFMs, coupled with the ambiguity and context-dependence of mobile tasks, makes LFM-based automation unreliable and prone to errors. To address this critical challenge, we introduce VeriSafe Agent (VSA) [^1] The system is available at: https://github.com/VeriSafeAgent/VeriSafeAgent: a formal verification system that serves as a logically grounded safeguard for Mobile GUI Agents. VSA deterministically ensures that an agent’s actions strictly align with user intent before executing the action. At its core, VSA introduces a novel autoformalization technique that translates natural language user instructions into a formally verifiable specification. This enables runtime, rule-based verification of agent’s actions, detecting erroneous actions even before they take effect. To the best of our knowledge, VSA is the first attempt to bring the rigor of formal verification to GUI agents, bridging the gap between LFM-driven actions and formal software verification. We implement VSA using off-the-shelf LFM services (GPT-4o) and evaluate its performance on 300 user instructions across 18 widely used mobile apps. The results demonstrate that VSA achieves 94.33%–98.33% accuracy in verifying agent actions, outperforming existing LFM-based verification methods by 30.00%–16.33%, and increases the GUI agent’s task completion rate by 90%–130%.
GUI Agent, AI Safety; Formal Verification
journalyear: 2025copyright: ccconference: The 31st Annual International Conference on Mobile Computing and Networking; November 4–8, 2025; Hong Kong, Chinabooktitle: The 31st Annual International Conference on Mobile Computing and Networking (ACM MOBICOM ’25), November 4–8, 2025, Hong Kong, Chinadoi: 10.1145/3680207.3765248isbn: 979-8-4007-1129-9/2025/11ccs: Computing methodologies Mobile agents
@@ 1. Introduction
However, unlike conventional software verification (Liu et al., 2023; Yang et al., 2024; Wu et al., 2024; Obi et al., 2025), which operates on predefined safety requirements, VSA faces the unique challenge of formalizing impromptu, user-defined instructions on-the-fly, across diverse and dynamic landscape of mobile applications. To accomplish this, VSA introduces three key innovations:n
@@
    Domain-Specific Language (DSL) and Developer Library [^1] The library is available at: https://github.com/VeriSafeAgent/VeriSafeAgent_Library: A specialized language and accompanying tools tailored to dynamic nature of mobile environments. Collectively, they can encode both the natural-language user instructions and corresponding UI actions as logical formulas.
@@
We implement a prototype of VSA using off-the-shelf online LFM service (GPT-4o) and integrate it with M3A (Multimodal Autonomous Agent for Android) GUI agent (Rawles et al., 2024). Our evaluation, conducted on 300 user instructions across 18 widely used mobile applications demonstrates that with an estimated $\sim$437 additional lines of code (LoC) per app, VSA successfully verifies GUI agent actions with up to 98.33% accuracy, achieving a false positive rate of 1.3% and a false negative rate of 0.3%. This significantly outperforms baseline reflection agents by 23.2%. Furthermore, VSA’s structured feedback, derived from logical verification results, enables the GUI agent to enhance its task completion rate by 130%, a substantial improvement compared to reflection agents, which failed to correct any tasks.
@@ 3.1. Challenges
C3. Simply flagging an action as incorrect is insufficient for improving the overall practicality of Mobile GUI Agents. To enhance their accuracy and reliability, we must provide actionable feedback that correctly guides agents toward the goal. However, LFM-generated feedback is often vague, underspecified, or even incorrect, potentially leading to repeated errors or suboptimal actions. Therefore, VSA incorporates a structured feedback generation mechanism (Figure 4) that systematically generates precise, rule-based feedback by identifying unmet or violated conditions based on the logical verification results.
@@ 3.2. System Workflow
App Development Phase. To enable rigorous rule-based verification, VSA requires app developers to define the critical states and transitions relevant for verification using VSA developer library (§ 4.2). This includes declaring application state spaces and their update mechanisms. While this requires additional developer effort, typically involving 5 to 10 lines of code per state, and potentially more depending on the application’s complexity, such practices are already common in modern app development workflows (Meta, 2025; Google, 2025; Apple, 2025).
@@
$\textsf{MovieInfo(title=``M'',\,time=7pm)}\rightarrow\textsf{Book}$
@@ 4.1. Design of Domain-Specific Language
{bnf}
@@
Constraint. A constraint $c$ has the form varopconstvar\;op\;const (e.g., $x=10$), where $x$ is a variable representing a specific element within the application state (e.g., restaurant_name, location). $op$ is a comparison operator, and constconst represents a constant value. The expression $x;op;const$ returns true if the variable $x$ meets the condition defined by the operator $op$ and the constant constconst. If $x$ has not been observed in the app, it defaults to undefined, making the constraint always false. The variable $x$ can be one of the following six types: string, number, Boolean, enumeration, date, and time. For example, given a constraint $x\geq 10$, if the value of the variable $x$ was observed to be 15 in the application, then the constraint is considered true.
@@
$R_{1}:$ $\textsf{RestaurantInfo}(\textsf{name}=\textsf{``R''})$
$\land\textsf{ReserveInfo}(\textsf{date}=\textsf{Today},\textsf{time}<\textsf{19:00},\textsf{available}=\textsf{True})$
$\rightarrow\textsf{Reserve}$
@@
$R_{3}:$ $\textsf{RestaurantInfo}(\textsf{name}=\textsf{``R''})$
$\land\textsf{ReserveInfo}(\textsf{date}=\textsf{Today},\textsf{time}<\textsf{19:00},\textsf{available}\neq\textsf{True})$
$\rightarrow\textsf{Done}$
@@ 4.2. Developer Library
To this end, VSA provides a developer library that empowers app developers to explicitly define application states (as state predicates) and their corresponding update mechanisms (as state transitions). This approach capitalizes on developers’ deep understanding of their application’s internal logic and intended behavior. Although it requires additional developer effort, it enables highly accurate state definitions suitable for rule-based logical verification—a level of precision that is unattainable with automated approaches like static code analysis or LFM-driven dynamic analysis.
@@
vsaStateManager.updateState("RestaurantInfo", {
@@
Another key advantage of developer-defined state predicates is the control over granularity. In many mobile tasks, individual atomic actions are not inherently meaningful. For example, in an e-commerce application, entering individual fields of a shipping address (street, city, zip code) is a set of preparatory steps. The critical state change is the submission of the complete address. Verifying every single atomic action would be inefficient and could mislead LFMs to think that some steps are incomplete. By allowing developers to strategically define verification checkpoints that have a meaningful impact on the task’s execution, VSA can significantly improve the efficiency and accuracy of task verification.
4.2.3. Post-action State Update Triggers.
While pre-action verification is ideal for preventing errors, some action outcomes are inherently unpredictable (e.g., the result of a network request). For these scenarios, VSA supports post-action state updates. Developers can call the ‘updateState()’ API directly, often within asynchronous callbacks, to update the state after the outcome of an action is known. This allows developers to handle dynamic application behaviors and real-world uncertainties such as network delays, asynchronous operations, error conditions, and dynamically determined content.
For instance, consider a flight booking app where the price and availability of flights change dynamically based on real-time data. A developer can define a FlightSearchResult state with variables for price, availability, and departure time. They can then associate a state update trigger with the asynchronous network callback that gets executed when the user clicks the “Search” button. Upon receiving the network response, the developer-defined trigger logic updates the FlightSearchResult based on the actual results. If the network fails or no seats are available, the trigger can handle these conditions appropriately, updating the state to reflect the error or unavailability. This ability to embed complex, conditional logic directly into the verification framework enables VSA to handle a wide range of user requirements and edge cases.
@@ 5.1. Intent Encoder
To ensure the encoded specification is correct and includes all relevant constraints, VSA employs two complementary self-corrective techniques: (i) rule-based syntax checking, and (ii) decoding-based semantics checking. These methods allow the LFM to incrementally identify, correct, and refine its translations, improving accuracy and reliability.
As an initial step, the Intent Encoder prompts the LFM with: i) the user’s natural language instruction and ii) the set of developer-defined states (from the developer library). From this information and user instruction, the Intent Encoder generates a draft specification representing the user’s intent using available states. The draft specification then undergoes two checking stages:
Rule-based Syntax Checking. Inspired by conventional programming languages, VSA checks that the generated Specification strictly adheres to the DSL’s syntax (§ 4.1.1) and that the types of constants $c$ used within the Specification are consistent with the types of variables declared in the developer-defined states. For example, if the LFM generates a state predicate like $\textsf{RestaurantInfo}(\textsf{name}\geq\textsf{100})$ while the type of the name variable is defined as ‘String’ in the developer library, this inconsistency is immediately flagged as a syntax error. If an error is detected, a structured error message is returned to the encoding LFM as feedback, guiding it toward a corrected translation.
Decoding-Based Semantics Checking. To further validate the semantic correctness of the translation, the generated Specification is decoded back into a natural language description using a separate decoder LFM. A checker LFM then compares this decoded description to the original user instruction, confirming that the encoding genuinely captures the user’s intent rather than merely manipulating symbols. If discrepancies are found, the checker LFM identifies the root cause and provides feedback to the encoding LFM.
These complementary checking mechanisms significantly improve the robustness and accuracy of the intent encoding process, effectively mitigating risks associated with solely relying on probabilistic LFM outputs.
@@
To address this challenge, VSA introduces an Experience-driven Encoding, which leverages previously successful encodings to guide subsequent translations. When an instruction is successfully translated and verified, its result specification is cached. Since a single specification consists of multiple rules representing an instruction’s intermediate objectives, numerous instructions for the same application may share similar rules. By retrieving and prioritizing predicates based on past encoding experience, we can significantly reduce the search space for the LFM encoder, improving both efficiency and consistency.
@@ 5.2. Intent Verifier
Predicate-level Verification is performed at every state update. Whenever a state predicate within the Specification is re-evaluated due to a variable update, its truth value is checked. If a predicate evaluates to false, this indicates a potential violation of the user’s intent. For example, if the Specification includes the predicate $\textsf{RestaurantInfo}(\textsf{name}=\textsf{``R''})$, and the agent’s action causes the name variable to be updated to “S”, the predicate would evaluate to false, indicating a potential violation of user intent.
@@
Rule-level Verification is performed at irreversible or critical checkpoints within the task execution. To identify these checkpoints, we modified the GUI agent’s prompt so that it generates a critical flag along with the action if that action is intended to directly achieve one of the objectives $o$ within the Specification (e.g., Reserve).
@@
This two-tiered verification approach guides the agent toward its objectives while ensuring the safe execution of high-risk actions.
@@ 7. Implementation
Figure 5. Verification accuracy
Our VSA implementation is designed as a modular component that can be seamlessly integrated with existing GUI agents without requiring modifications to their underlying architecture. VSA is implemented using Python, and its Intent Encoder leverages off-the-shelf LFM GPT-4o.
Developer Library Guidelines. Developers should carefully determine the granularity of State Predicates to suit their application context. Overly fine-grained State Predicates can make the encoded specification too detailed, increasing the false positive rate. Conversely, coarse-grained State Predicates may introduce loopholes in the encoded specification, leading to a higher false negative rate. To assist with finding the sweet spot, the VSA Library provides detailed documentation and guidelines on how to effectively define State Predicates. Given the potential severity of false negatives, it’s recommended to begin with a fine-grained verification first and then iteratively refine the predicates to minimize false positives.
String Similarity. To compare strings semantically rather than structurally, we use OpenAI’s text-embedding-3-small model (openai, 2023) and compute cosine similarity. Based on empirical studies, we set a similarity threshold of 0.7 to determine equivalence.
@@ 8. Evaluation
In this section, we evaluate the performance of VSA against baseline guardrail frameworks in terms of i) verification accuracy, ii) feedback effect, and iii) latency and cost. Throughout the experiments, we integrated each framework, including VSAwith the M3A mobile GUI agent (Rawles et al., 2024)—a simple yet powerful mobile GUI agent powered by GPT-4o—and conducted evaluations using a Google Pixel 8 smartphone.
Baselines & VSA Variants. To evaluate VSA, we compare it with two LFM-based reflection schemes: Pre-action reflection (Li et al., 2024a) and Post-action reflection (Wang et al., 2024), both of which utilize GPT-4o as their backbone LFM. For VSA, we evaluate two variants: VSA-cold and VSA-warm. VSA-cold encodes user instructions solely using self-corrective encoding, without cached memory. On the other hand, VSA-warm leverages both self-corrective encoding and experience-driven encoding, assuming that a cached list of candidate predicates is available for the given instruction (see § 5.1 for details).
@@ 8.1. Dataset
To comprehensively evaluate VSA, we constructed a dataset [^1] The dataset is available at: https://github.com/VeriSafeAgent/VeriSafeAgent of 300 user instructions, divided into two complementary subsets of 150 instructions each: the Correct dataset and the Wrong dataset.
Each instruction is paired with an execution path—a sequence of UI actions performed by the GUI agent. In the Correct dataset, instructions are matched with execution paths that accurately fulfill the user’s intent. In contrast, the Wrong dataset deliberately introduces mismatches between instructions and their corresponding execution paths. Specifically, we altered key details in the original Correct instructions to create discrepancies. For example, ”On Play Books, turn to the search page and check for top-selling books” was modified to ”On Play Books, turn to the search page and check for newly released books”, making the original execution path no longer satisfy the updated instruction. Such cases test the verification framework’s ability to detect errors, enabling evaluation of both True Positives (correctly rejecting invalid executions) and False Negatives (incorrectly accepting invalid executions).
Of the 150 unique instructions, 125 were sampled from the LlamaTouch dataset (Zhang et al., 2024), selected based on app popularity (global usage), instruction realism (likelihood of real-world occurrence), and current app availability. We excluded tasks that were overly localized, deprecated, or overly specific and prone to overfitting. The remaining 25 instructions were newly created to form the Challenge dataset, designed to push the system to its limits with longer action sequences, multiple constraints, and conditional branches. Spanning 9 different applications, these instructions average 19.16 steps (maximum 40), far exceeding LlamaTouch’s average of 5.67 steps.
Finally, as described in § 4.2, VSA relies on developer-defined state specifications and state-update triggers. To emulate this in closed-source mobile apps, we manually annotated state predicates and state transitions for each GUI screen and action in the dataset. This manual annotation serves as a surrogate for direct integration with the developer library, effectively replicating the library’s intended functionality.